# PICore: Physics-Informed Unsupervised Coreset Selection for Data Efficient Neural Operator Training

### v1

@@ A.3 Darcy Flow
We obtain the steady-state solution of Darcy’s equation on the unit square by evolving a time-dependent problem until convergence. The target elliptic problem is
@@
with an appropriate random-field initial condition and homogeneous Dirichlet boundary data. We use the strong form $\nabla\cdot(a\nabla u)-f$ for the residual as in Li et al. (2024c).
A.4 Navier-Stoker Equation
@@
[…]
@@
The solution is obtained on a $256\times 256$ grid via a Fourier pseudospectral scheme: first, we solve $\Delta\psi=-\omega$ in Fourier space to recover the stream function $\psi$ and velocity $u$, then compute the nonlinear advection term $u\cdot\nabla\omega$ in physical space with a $2/3$-dealiasing filter, and finally advance in time using Crank–Nicolson for diffusion coupled with an explicit update for the nonlinear term.
@@ Appendix B Coreset Selection Algorithms
$A^{*}=\operatorname*{arg\,min}_{A\subset V}|S|,\;\sum_{n\in V}\min_{m\in S}\max_{\theta}||g_{n}-g_{m}||$
so every $g_{i}$ is “covered” by its most similar selected gradient. CRAIG selects the smallest subset S such that every example gradient is close (in $\mathcal{L}_{2}$) to at least one gradient in S. We approximate the coverage objective with the stochastic-greedy algorithm applied to the pairwise Euclidean similarity matrix of last-layer gradients. Greedy (or stochastic-greedy) selection gives a $(1-1/e)$-approximation in finite similarity evaluations. After $S$ is chosen, CRAIG sets integer weights
@@ Appendix C Further Results
C.1 Zero Shot Super Resolution
@@
Since Neural Operators learn parameters independently of the discretization (unlike PINNs), trained neural operators can perform zero-shot super-resolution, which allows for training a model at a lower resolution and evaluating at a higher resolution. We scale the Advection and Burgers datasets to resolutions of 128 and 256, the Darcy dataset to 128, and the Navier Stokes Incompressible dataset to size 256 in tables  914.
@@
gradmatch $8.38\times 10^{-2}$ $7.87\times 10^{-2}$ $7.85\times 10^{-2}$ $7.86\times 10^{-2}$ $7.73\times 10^{-2}$ $5.90\times 10^{-2}$
@@
graNd $8.38\times 10^{-2}$ $8.15\times 10^{-2}$ $7.86\times 10^{-2}$ $7.72\times 10^{-2}$ $7.50\times 10^{-2}$ $5.90\times 10^{-2}$
UNO Supervised craig $1.72\times 10^{-1}$ $1.66\times 10^{-1}$ $1.61\times 10^{-1}$ $1.48\times 10^{-1}$ $1.14\times 10^{-1}$ $1.09\times 10^{-1}$
@@
gradmatch $9.22\times 10^{-2}$ $8.79\times 10^{-2}$ $8.80\times 10^{-2}$ $8.85\times 10^{-2}$ $8.76\times 10^{-2}$ $7.07\times 10^{-2}$
adacore $9.07\times 10^{-2}$ $8.65\times 10^{-2}$ $8.87\times 10^{-2}$ $8.43\times 10^{-2}$ $8.47\times 10^{-2}$ $7.07\times 10^{-2}$
@@
gradmatch $1.84\times 10^{-1}$ $1.79\times 10^{-1}$ $1.77\times 10^{-1}$ $1.59\times 10^{-1}$ $1.41\times 10^{-1}$ $1.31\times 10^{-1}$
@@
graNd $1.83\times 10^{-1}$ $1.79\times 10^{-1}$ $1.77\times 10^{-1}$ $1.68\times 10^{-1}$ $1.43\times 10^{-1}$ $1.31\times 10^{-1}$
@@
gradmatch $6.56\times 10^{-2}$ $6.35\times 10^{-2}$ $6.68\times 10^{-2}$ $6.32\times 10^{-2}$ $6.36\times 10^{-2}$ $4.57\times 10^{-2}$
@@
gradmatch $6.40\times 10^{-2}$ $6.62\times 10^{-2}$ $6.81\times 10^{-2}$ $6.47\times 10^{-2}$ $6.15\times 10^{-2}$ $4.57\times 10^{-2}$
@@
el2n $6.29\times 10^{-2}$ $6.36\times 10^{-2}$ $6.60\times 10^{-2}$ $6.29\times 10^{-2}$ $6.38\times 10^{-2}$ $4.57\times 10^{-2}$
@@
adacore $4.36\times 10^{-2}$ $3.39\times 10^{-2}$ $2.98\times 10^{-2}$ $2.47\times 10^{-2}$ $2.23\times 10^{-2}$ $2.16\times 10^{-2}$
el2n $2.94\times 10^{-2}$ $2.76\times 10^{-2}$ $2.62\times 10^{-2}$ $2.39\times 10^{-2}$ $2.28\times 10^{-2}$ $2.16\times 10^{-2}$
@@
gradmatch $2.97\times 10^{-2}$ $2.76\times 10^{-2}$ $2.69\times 10^{-2}$ $2.40\times 10^{-2}$ $2.27\times 10^{-2}$ $2.16\times 10^{-2}$
@@
el2n $2.96\times 10^{-2}$ $2.77\times 10^{-2}$ $2.62\times 10^{-2}$ $2.43\times 10^{-2}$ $2.29\times 10^{-2}$ $2.16\times 10^{-2}$
graNd $2.89\times 10^{-2}$ $2.72\times 10^{-2}$ $2.66\times 10^{-2}$ $2.49\times 10^{-2}$ $2.29\times 10^{-2}$ $2.16\times 10^{-2}$
@@
gradmatch $6.72\times 10^{-2}$ $6.55\times 10^{-2}$ $6.89\times 10^{-2}$ $6.55\times 10^{-2}$ $6.60\times 10^{-2}$ $4.89\times 10^{-2}$
@@
el2n $7.04\times 10^{-2}$ $7.38\times 10^{-2}$ $6.96\times 10^{-2}$ $6.66\times 10^{-2}$ $6.32\times 10^{-2}$ $4.89\times 10^{-2}$
graNd $6.89\times 10^{-2}$ $6.21\times 10^{-2}$ $6.69\times 10^{-2}$ $6.50\times 10^{-2}$ $6.29\times 10^{-2}$ $4.89\times 10^{-2}$
PICore craig $6.85\times 10^{-2}$ $6.52\times 10^{-2}$ $6.59\times 10^{-2}$ $6.45\times 10^{-2}$ $6.63\times 10^{-2}$ $4.89\times 10^{-2}$
@@
adacore $7.50\times 10^{-2}$ $6.75\times 10^{-2}$ $6.80\times 10^{-2}$ $6.61\times 10^{-2}$ $6.69\times 10^{-2}$ $4.89\times 10^{-2}$
el2n $6.49\times 10^{-2}$ $6.57\times 10^{-2}$ $6.82\times 10^{-2}$ $6.52\times 10^{-2}$ $6.62\times 10^{-2}$ $4.89\times 10^{-2}$
graNd $6.88\times 10^{-2}$ $6.78\times 10^{-2}$ $7.12\times 10^{-2}$ $6.72\times 10^{-2}$ $6.69\times 10^{-2}$ $4.89\times 10^{-2}$
UNO Supervised craig $3.38\times 10^{-2}$ $3.36\times 10^{-2}$ $3.26\times 10^{-2}$ $3.06\times 10^{-2}$ $2.96\times 10^{-2}$ $2.92\times 10^{-2}$
gradmatch $3.52\times 10^{-2}$ $3.43\times 10^{-2}$ $3.36\times 10^{-2}$ $3.22\times 10^{-2}$ $3.10\times 10^{-2}$ $2.92\times 10^{-2}$
@@
el2n $3.46\times 10^{-2}$ $3.33\times 10^{-2}$ $3.27\times 10^{-2}$ $3.11\times 10^{-2}$ $3.05\times 10^{-2}$ $2.92\times 10^{-2}$
graNd $3.51\times 10^{-2}$ $3.40\times 10^{-2}$ $3.34\times 10^{-2}$ $3.22\times 10^{-2}$ $3.11\times 10^{-2}$ $2.92\times 10^{-2}$
PICore craig $3.63\times 10^{-2}$ $3.47\times 10^{-2}$ $3.36\times 10^{-2}$ $3.11\times 10^{-2}$ $2.97\times 10^{-2}$ $2.92\times 10^{-2}$
gradmatch $3.50\times 10^{-2}$ $3.35\times 10^{-2}$ $3.34\times 10^{-2}$ $3.14\times 10^{-2}$ $3.08\times 10^{-2}$ $2.92\times 10^{-2}$
@@
el2n $3.47\times 10^{-2}$ $3.36\times 10^{-2}$ $3.29\times 10^{-2}$ $3.16\times 10^{-2}$ $3.09\times 10^{-2}$ $2.92\times 10^{-2}$
graNd $3.41\times 10^{-2}$ $3.34\times 10^{-2}$ $3.34\times 10^{-2}$ $3.28\times 10^{-2}$ $3.12\times 10^{-2}$ $2.92\times 10^{-2}$
@@
FNO Supervised craig $1.67\times 10^{-1}$ $1.51\times 10^{-1}$ $1.36\times 10^{-1}$ $1.15\times 10^{-1}$ $1.02\times 10^{-1}$ $8.76\times 10^{-2}$
gradmatch $1.72\times 10^{-1}$ $1.47\times 10^{-1}$ $1.36\times 10^{-1}$ $1.15\times 10^{-1}$ $1.04\times 10^{-1}$ $8.76\times 10^{-2}$
adacore $1.88\times 10^{-1}$ $1.61\times 10^{-1}$ $1.41\times 10^{-1}$ $1.18\times 10^{-1}$ $1.05\times 10^{-1}$ $8.76\times 10^{-2}$
el2n $1.68\times 10^{-1}$ $1.42\times 10^{-1}$ $1.33\times 10^{-1}$ $1.15\times 10^{-1}$ $1.02\times 10^{-1}$ $8.76\times 10^{-2}$
graNd $1.67\times 10^{-1}$ $1.46\times 10^{-1}$ $1.32\times 10^{-1}$ $1.16\times 10^{-1}$ $1.04\times 10^{-1}$ $8.76\times 10^{-2}$
PICore craig $1.68\times 10^{-1}$ $1.53\times 10^{-1}$ $1.39\times 10^{-1}$ $1.16\times 10^{-1}$ $1.03\times 10^{-1}$ $8.76\times 10^{-2}$
gradmatch $1.73\times 10^{-1}$ $1.48\times 10^{-1}$ $1.38\times 10^{-1}$ $1.16\times 10^{-1}$ $1.03\times 10^{-1}$ $8.76\times 10^{-2}$
adacore $1.98\times 10^{-1}$ $1.58\times 10^{-1}$ $1.37\times 10^{-1}$ $1.10\times 10^{-1}$ $1.02\times 10^{-1}$ $8.76\times 10^{-2}$
el2n $1.66\times 10^{-1}$ $1.48\times 10^{-1}$ $1.32\times 10^{-1}$ $1.13\times 10^{-1}$ $1.03\times 10^{-1}$ $8.76\times 10^{-2}$
graNd $1.72\times 10^{-1}$ $1.47\times 10^{-1}$ $1.35\times 10^{-1}$ $1.13\times 10^{-1}$ $1.03\times 10^{-1}$ $8.76\times 10^{-2}$

### v2

@@ A.3 Darcy Flow
We obtain the steady‐state solution of Darcy’s equation on the unit square by evolving a time‐dependent problem until convergence. The target elliptic problem is
@@
with an appropriate random‐field initial condition and homogeneous Dirichlet boundary data. We use the strong form $\nabla\cdot(a\nabla u)-f$ for the residual as in Li et al. (2024c).
A.4 Navier-Stokes Equation
@@
$\partial_{t}\omega+u\cdot\nabla\omega=\nu\,\Delta\omega+f,\quad\nabla\!\cdot u=0,\quad\omega(x,0)\sim\mathcal{N}\bigl(0,\;7^{3/2}(-\Delta+49I)^{-2.5}\bigr),$
@@
The solution is obtained on a $256\times 256$ grid via a Fourier pseudospectral scheme: first, we solve $\Delta\psi=-\omega$ in Fourier space to recover the stream function $\psi$ and velocity $u$, then compute the nonlinear advection term $u\cdot\nabla\omega$ in physical space with a 2/32/3‐dealiasing filter, and finally advance in time using Crank–Nicolson for diffusion coupled with an explicit update for the nonlinear term.
@@ Appendix B Coreset Selection Algorithms
[…]
so every $g_{i}$ is “covered” by its most similar selected gradient. CRAIG selects the smallest subset S such that every example gradient is close (in $\mathcal{L}_{2}$) to at least one gradient in S. We approximate the coverage objective with the stochastic-greedy algorithm applied to the pairwise Euclidean similarity matrix of last-layer gradients. Greedy (or stochastic-greedy) selection gives a (11/e)(1-1/e)-approximation in finite similarity evaluations. After $S$ is chosen, CRAIG sets integer weights
@@ Appendix C Further Results
C.1 Unsupervised Selection Methods Comparison
Figure 4: Test NRMSE on the Advection dataset at resolution 64 across varying coreset percentages (20%–100%) between unsupervised and PICore-based coreset selection methods using both FNO and UNO architectures.
Figure 5: Test NRMSE on the Burgers dataset at resolution 64 across varying coreset percentages (20%–100%) between unsupervised and PICore-based coreset selection methods using both FNO and UNO architectures.
Figure 6: Test NRMSE on the Darcy dataset at resolution 64 across varying coreset percentages (20%–100%) between unsupervised and PICore-based coreset selection methods using both FNO and UNO architectures.
Figure 7: Test NRMSE on the Navier Stokes Incompressible dataset at resolution 64 across varying coreset percentages (20%–100%) between unsupervised and PICore-based coreset selection methods using both FNO and UNO architectures.
C.2 Spatial Orientation of Supervised Coreset Selection and PICore
To better understand the differences between supervised coreset selection and PICore, we analyze how well each method covers the input space by computing the average distance from coreset points to their centroid, which serves as a proxy for spread or diversity. We compute this distance with respect to the $\|\cdot\|_{L^{2}(\Omega)}$ norm, where the centroid is the average data point element-wise and the average distance is the average norm between the centroid and the selected data points in the coreset. As shown in Figure 8, this distance is nearly identical across datasets and neural operators (FNO and UNO), with overlapping standard error bars with differences decreasing as the PDE complexity increases (Advection to Navier Stokes). This suggests that PICore selects coresets that are as well-distributed as those from supervised methods, despite not using labeled data. The comparable coverage indicates that differences in downstream performance likely arise from the type of points selected rather than their spatial distribution.
(a) Advection
(b) Burgers
(c) Darcy
(d) Navier Stokes
Figure 8: Average centroid distances across datasets for FNO and UNO.
C.3 Ablation Study
Table 9: Ablation on Warm Start at resolution 64
Operator $T_{w}$ 20.0% 30.0% 40.0% 60.0% 80.0% 100.0%
FNO 10 epochs $64\pm 0.14\times 10^{-2} 1.1$$14\pm 0.10\times 10^{-2}$ $8.07\pm 0.48\times 10^{-3}$ $5.32\pm 0.30\times 10^{-3}$ $4.17\pm 0.17\times 10^{-3}$ $3.95\pm 0.10\times 10^{-3}$
$(3.23\times)$ $(2.48\times)$ $(2.06\times)$ $(1.52\times)$ $(1.19\times)$ $(1.00\times)$
25 epochs $1.52\pm 0.13\times 10^{-2}$ $1.07\pm 0.06\times 10^{-2}$ $7.71\pm 0.33\times 10^{-3}$ $5.33\pm 0.14\times 10^{-3}$ $4.15\pm 0.06\times 10^{-3}$ $3.95\pm 0.10\times 10^{-3}$
$(3.18\times)$ $(2.44\times)$ $(2.02\times)$ $(1.48\times)$ $(1.17\times)$ $(1.00\times)$
50 epochs $69\pm 0.07\times 10^{-2} 1.1$$12\pm 0.08\times 10^{-2}$ $7.96\pm 0.38\times 10^{-3}$ $5.51\pm 0.17\times 10^{-3}$ $4.16\pm 0.11\times 10^{-3}$ $3.95\pm 0.10\times 10^{-3}$
$(3.10\times)$ $(2.36\times)$ $(1.95\times)$ $(1.43\times)$ $(1.13\times)$ $(1.00\times)$
100 epochs $1.69\pm 0.14\times 10^{-2}$ $1.06\pm 0.04\times 10^{-2}$ $8.24\pm 0.45\times 10^{-3}$ $5.40\pm 0.16\times 10^{-3}$ $4.15\pm 0.05\times 10^{-3}$ $3.95\pm 0.10\times 10^{-3}$
$(2.94\times)$ $(2.21\times)$ $(1.83\times)$ $(1.34\times)$ $(1.05\times)$ $(1.00\times)$
UNO 10 epochs $2.97\pm 0.04\times 10^{-2}$ $2.44\pm 0.06\times 10^{-2}$ $2.13\pm 0.04\times 10^{-2}$ $1.77\pm 0.04\times 10^{-2}$ $1.57\pm 0.01\times 10^{-2}$ $1.49\pm 0.04\times 10^{-2}$
$(3.79\times)$ $(2.79\times)$ $(2.25\times)$ $(1.60\times)$ $(1.24\times)$ $(1.00\times)$
25 epochs $2.99\pm 0.04\times 10^{-2}$ $2.51\pm 0.04\times 10^{-2}$ $2.09\pm 0.02\times 10^{-2}$ $1.76\pm 0.03\times 10^{-2}$ $1.57\pm 0.04\times 10^{-2}$ $1.49\pm 0.04\times 10^{-2}$
$(3.72\times)$ $(2.73\times)$ $(2.19\times)$ $(1.56\times)$ $(1.20\times)$ $(1.00\times)$
50 epochs $2.89\pm 0.03\times 10^{-2}$ $2.42\pm 0.03\times 10^{-2}$ $2.18\pm 0.05\times 10^{-2}$ $1.73\pm 0.03\times 10^{-2}$ $1.54\pm 0.02\times 10^{-2}$ $1.49\pm 0.04\times 10^{-2}$
$(3.60\times)$ $(2.64\times)$ $(2.11\times)$ $(1.49\times)$ $(1.15\times)$ $(1.00\times)$
100 epochs $3.00\pm 0.08\times 10^{-2}$ $2.53\pm 0.05\times 10^{-2}$ $2.05\pm 0.04\times 10^{-2}$ $1.73\pm 0.03\times 10^{-2}$ $1.55\pm 0.02\times 10^{-2}$ $1.49\pm 0.04\times 10^{-2}$
$(3.38\times)$ $(2.46\times)$ $(1.96\times)$ $(1.38\times)$ $(1.07\times)$ $(1.00\times)$
The ablation study in Table 9 examines how different warm start epochs influence the performance of PICore on FNO and UNO. We fix the coreset selection algorithm to EL2N and the dataset to Burgers. Across warm start configurations, extending the number of epochs beyond 10 to 25 or 50 epochs leads to only marginal gains, and by 100 epochs, the improvements are negligible or even slightly inconsistent within the bounds of standard deviation. For FNO, the 25-epoch variant achieves the lowest errors at lower data fractions (20–40%), suggesting that moderate warm starting may yield slightly better initialization for PICore. UNO, on the other hand, shows stable performance across all pretraining lengths, implying that it benefits less from extended warm starting. Overall, the table suggests that minimal warm starting (around 10–25 epochs) is sufficient for both operator families, while additional finetuning offers little benefit.
C.4 Multi-Resolution Coreset Selection
Table 10: Burgers NRMSE with Multi Resolution Data
Operator Method Algorithm 20.0% 30.0% 40.0% 60.0% 80.0% 100.0%
FNO Supervised craig $8.29\pm 1.58\times 10^{-2}$ $7.73\pm 0.83\times 10^{-2}$ $56\pm 0.75\times 10^{-2} 7.7$$78\pm 0.88\times 10^{-2}$ $7.25\pm 0.84\times 10^{-2}$ $6.85\pm 0.79\times 10^{-2}$
gradmatch $8.33\pm 1.16\times 10^{-2}$ $8.10\pm 0.93\times 10^{-2}$ $7.92\pm 0.44\times 10^{-2}$ $7.64\pm 0.81\times 10^{-2}$ $6.94\pm 0.58\times 10^{-2}$ $6.85\pm 0.79\times 10^{-2}$
adacore $9.24\pm 0.83\times 10^{-2}$ $8.73\pm 0.54\times 10^{-2}$ $8.03\pm 0.98\times 10^{-2}$ $7.44\pm 0.93\times 10^{-2}$ $6.89\pm 0.75\times 10^{-2}$ $6.85\pm 0.79\times 10^{-2}$
el2n $7.61\pm 0.96\times 10^{-2}$ $\mathbf{7.34\pm 1.30\times 10^{-2}}$ $7.55\pm 0.60\times 10^{-2}$ $7.30\pm 0.61\times 10^{-2}$ $6.94\pm 0.85\times 10^{-2}$ $6.85\pm 0.79\times 10^{-2}$
graNd $8.44\pm 0.73\times 10^{-2}$ $95\pm 1.00\times 10^{-2} 7.7$$75\pm 0.88\times 10^{-2}$ $7.45\pm 0.60\times 10^{-2}$ $7.23\pm 0.58\times 10^{-2}$ $6.85\pm 0.79\times 10^{-2}$
Acceleration $3.75\pm 0.00\times$ $76\pm 0.00\times 2.2$$24\pm 0.00\times$ $1.60\pm 0.00\times$ $1.24\pm 0.00\times$ $1.00\pm 0.00\times$
PICore craig $\mathbf{7.37\pm 1.63\times 10^{-2}}$ $8.09\pm 0.62\times 10^{-2}$ $7.70\pm 1.20\times 10^{-2}$ $7.36\pm 0.70\times 10^{-2}$ $7.31\pm 0.45\times 10^{-2}$ $6.85\pm 0.79\times 10^{-2}$
gradmatch $7.90\pm 1.89\times 10^{-2}$ $8.07\pm 1.55\times 10^{-2}$ $7.79\pm 1.19\times 10^{-2}$ $7.54\pm 0.59\times 10^{-2}$ $7.15\pm 0.85\times 10^{-2}$ $6.85\pm 0.79\times 10^{-2}$
adacore $8.56\pm 1.78\times 10^{-2}$ $8.14\pm 1.36\times 10^{-2}$ $8.20\pm 0.81\times 10^{-2}$ $7.90\pm 1.03\times 10^{-2}$ $7.34\pm 1.10\times 10^{-2}$ $6.85\pm 0.79\times 10^{-2}$
el2n $8.10\pm 0.66\times 10^{-2}$ $7.37\pm 1.87\times 10^{-2}$ $\mathbf{7.49\pm 1.54\times 10^{-2}}$ $\mathbf{7.29\pm 0.65\times 10^{-2}}$ $\mathbf{6.78\pm 0.73\times 10^{-2}}$ $6.85\pm 0.79\times 10^{-2}$
graNd $8.88\pm 0.63\times 10^{-2}$ $8.39\pm 0.87\times 10^{-2}$ $8.19\pm 1.16\times 10^{-2}$ $7.56\pm 0.61\times 10^{-2}$ $7.31\pm 0.81\times 10^{-2}$ $6.85\pm 0.79\times 10^{-2}$
Acceleration $5.08\pm 0.01\times$ $3.32\pm 0.00\times$ $2.54\pm 0.00\times$ $1.70\pm 0.00\times$ $1.27\pm 0.00\times$ $1.00\pm 0.00\times$
Table 10 compares Supervised and Physics-Informed Coreset Selection (PICore) across varying data resolutions for the Burgers’ equation using the FNO operator. In this experiment, the dataset is split into two equal parts: one half is generated at a resolution of 32, and the other at a resolution of 64, introducing multi-scale variability in the training data. Overall, PICore demonstrates improved stability and often superior accuracy at lower data percentages, particularly in the FNO setting. For instance, under FNO with 20–60% data, PICore achieves consistently lower NRMSE values than supervised methods, especially with the el2n and craig algorithms, indicating better generalization when fewer samples are available. This advantage reflects the integration of physical constraints, which act as a strong inductive bias that helps retain key solution structures even when data is sparse or noisy. The mixed-resolution setup further amplifies this effect, as while supervised coresets can struggle to reconcile the differences between coarse (32) and fine (64) grids, PICore leverages physics-informed consistency to align features across scales, resulting in more stable and resolution-invariant representations of the Burgers’ dynamics. However, this setup also introduces greater variability and noise, as discrepancies between resolutions can create inconsistencies in gradient magnitudes and feature smoothness, but these issues are generally inherent to multi-resolution datasets, rather than specific to the coreset selection approach itself.
C.5 Zero Shot Super Resolution
@@
Since Neural Operators learn parameters independently of the discretization (unlike PINNs), trained neural operators can perform zero-shot super-resolution, which allows for training a model at a lower resolution and evaluating at a higher resolution. We scale the Advection and Burgers datasets to resolutions of 128 and 256, the Darcy dataset to 128, and the Navier Stokes Incompressible dataset to size 256 in tables  1116.
@@
gradmatch $8.38\times 10^{-2}$ $7.87\times 10^{-2}$ $7.85\times 10^{-2}$ $86\times 10^{-2} 7.7$$73\times 10^{-2}$ $5.90\times 10^{-2}$
@@
graNd $8.38\times 10^{-2}$ $8.15\times 10^{-2}$ $86\times 10^{-2} 7.7$$72\times 10^{-2}$ $7.50\times 10^{-2}$ $5.90\times 10^{-2}$
UNO Supervised craig $1.72\times 10^{-1}$ $1.66\times 10^{-1}$ $1.61\times 10^{-1}$ $48\times 10^{-1} 1.1$$14\times 10^{-1}$ $1.09\times 10^{-1}$
@@
gradmatch $9.22\times 10^{-2}$ $79\times 10^{-2} 8.8$$0\times 10^{-2} 8.8$$85\times 10^{-2}$ $8.76\times 10^{-2}$ $7.07\times 10^{-2}$
adacore $9.07\times 10^{-2}$ $65\times 10^{-2} 8.8$$87\times 10^{-2}$ $8.43\times 10^{-2}$ $8.47\times 10^{-2}$ $7.07\times 10^{-2}$
@@
gradmatch $1.84\times 10^{-1}$ $9\times 10^{-1} 1.77×1$$7\times 10^{-1}$ $1.59\times 10^{-1}$ $1.41\times 10^{-1}$ $1.31\times 10^{-1}$
@@
graNd $1.83\times 10^{-1}$ $9\times 10^{-1} 1.77×1$$7\times 10^{-1}$ $1.68\times 10^{-1}$ $1.43\times 10^{-1}$ $1.31\times 10^{-1}$
@@
gradmatch $6.56\times 10^{-2}$ $35\times 10^{-2} 6.6$$68\times 10^{-2}$ $6.32\times 10^{-2}$ $6.36\times 10^{-2}$ $4.57\times 10^{-2}$
@@
gradmatch $40\times 10^{-2} 6.6$$62\times 10^{-2}$ $6.81\times 10^{-2}$ $6.47\times 10^{-2}$ $6.15\times 10^{-2}$ $4.57\times 10^{-2}$
@@
el2n $6.29\times 10^{-2}$ $36\times 10^{-2} 6.6$$60\times 10^{-2}$ $6.29\times 10^{-2}$ $6.38\times 10^{-2}$ $4.57\times 10^{-2}$
@@
adacore $4.36\times 10^{-2}$ $3.39\times 10^{-2}$ $2.98\times 10^{-2}$ $47\times 10^{-2} 2.2$$23\times 10^{-2}$ $2.16\times 10^{-2}$
el2n $2.94\times 10^{-2}$ $2.76\times 10^{-2}$ $2.62\times 10^{-2}$ $39\times 10^{-2} 2.2$$28\times 10^{-2}$ $2.16\times 10^{-2}$
@@
gradmatch $2.97\times 10^{-2}$ $2.76\times 10^{-2}$ $2.69\times 10^{-2}$ $40\times 10^{-2} 2.2$$27\times 10^{-2}$ $2.16\times 10^{-2}$
@@
el2n $2.96\times 10^{-2}$ $2.77\times 10^{-2}$ $2.62\times 10^{-2}$ $43\times 10^{-2} 2.2$$29\times 10^{-2}$ $2.16\times 10^{-2}$
graNd $2.89\times 10^{-2}$ $2.72\times 10^{-2}$ $2.66\times 10^{-2}$ $49\times 10^{-2} 2.2$$29\times 10^{-2}$ $2.16\times 10^{-2}$
@@
gradmatch $6.72\times 10^{-2}$ $6.55\times 10^{-2}$ $6.89\times 10^{-2}$ $55\times 10^{-2} 6.6$$60\times 10^{-2}$ $4.89\times 10^{-2}$
@@
el2n $7.04\times 10^{-2}$ $7.38\times 10^{-2}$ $96\times 10^{-2} 6.6$$66\times 10^{-2}$ $6.32\times 10^{-2}$ $4.89\times 10^{-2}$
graNd $6.89\times 10^{-2}$ $21\times 10^{-2} 6.6$$69\times 10^{-2}$ $6.50\times 10^{-2}$ $6.29\times 10^{-2}$ $4.89\times 10^{-2}$
PICore craig $6.85\times 10^{-2}$ $6.52\times 10^{-2}$ $6.59\times 10^{-2}$ $45\times 10^{-2} 6.6$$63\times 10^{-2}$ $4.89\times 10^{-2}$
@@
adacore $7.50\times 10^{-2}$ $6.75\times 10^{-2}$ $80\times 10^{-2} 6.6$$1\times 10^{-2} 6.6$$69\times 10^{-2}$ $4.89\times 10^{-2}$
el2n $6.49\times 10^{-2}$ $6.57\times 10^{-2}$ $6.82\times 10^{-2}$ $52\times 10^{-2} 6.6$$62\times 10^{-2}$ $4.89\times 10^{-2}$
graNd $6.88\times 10^{-2}$ $6.78\times 10^{-2}$ $7.12\times 10^{-2}$ $72\times 10^{-2} 6.6$$69\times 10^{-2}$ $4.89\times 10^{-2}$
UNO Supervised craig $38\times 10^{-2} 3.3$$36\times 10^{-2}$ $3.26\times 10^{-2}$ $3.06\times 10^{-2}$ $2.96\times 10^{-2}$ $2.92\times 10^{-2}$
gradmatch $3.52\times 10^{-2}$ $43\times 10^{-2} 3.3$$36\times 10^{-2}$ $3.22\times 10^{-2}$ $3.10\times 10^{-2}$ $2.92\times 10^{-2}$
@@
el2n $46\times 10^{-2} 3.3$$33\times 10^{-2}$ $3.27\times 10^{-2}$ $3.11\times 10^{-2}$ $3.05\times 10^{-2}$ $2.92\times 10^{-2}$
graNd $3.51\times 10^{-2}$ $40\times 10^{-2} 3.3$$34\times 10^{-2}$ $3.22\times 10^{-2}$ $3.11\times 10^{-2}$ $2.92\times 10^{-2}$
PICore craig $3.63\times 10^{-2}$ $47\times 10^{-2} 3.3$$36\times 10^{-2}$ $3.11\times 10^{-2}$ $2.97\times 10^{-2}$ $2.92\times 10^{-2}$
gradmatch $50\times 10^{-2} 3.3$$5\times 10^{-2} 3.3$$34\times 10^{-2}$ $3.14\times 10^{-2}$ $3.08\times 10^{-2}$ $2.92\times 10^{-2}$
@@
el2n $47\times 10^{-2} 3.3$$36\times 10^{-2}$ $3.29\times 10^{-2}$ $3.16\times 10^{-2}$ $3.09\times 10^{-2}$ $2.92\times 10^{-2}$
graNd $41\times 10^{-2} 3.3$$4\times 10^{-2} 3.3$$34\times 10^{-2}$ $3.28\times 10^{-2}$ $3.12\times 10^{-2}$ $2.92\times 10^{-2}$
@@
FNO Supervised craig $1.67\times 10^{-1}$ $1.51\times 10^{-1}$ $36\times 10^{-1} 1.1$$15\times 10^{-1}$ $1.02\times 10^{-1}$ $8.76\times 10^{-2}$
gradmatch $1.72\times 10^{-1}$ $1.47\times 10^{-1}$ $36\times 10^{-1} 1.1$$15\times 10^{-1}$ $1.04\times 10^{-1}$ $8.76\times 10^{-2}$
adacore $1.88\times 10^{-1}$ $1.61\times 10^{-1}$ $41\times 10^{-1} 1.1$$18\times 10^{-1}$ $1.05\times 10^{-1}$ $8.76\times 10^{-2}$
el2n $1.68\times 10^{-1}$ $1.42\times 10^{-1}$ $33\times 10^{-1} 1.1$$15\times 10^{-1}$ $1.02\times 10^{-1}$ $8.76\times 10^{-2}$
graNd $1.67\times 10^{-1}$ $1.46\times 10^{-1}$ $32\times 10^{-1} 1.1$$16\times 10^{-1}$ $1.04\times 10^{-1}$ $8.76\times 10^{-2}$
PICore craig $1.68\times 10^{-1}$ $1.53\times 10^{-1}$ $39\times 10^{-1} 1.1$$16\times 10^{-1}$ $1.03\times 10^{-1}$ $8.76\times 10^{-2}$
gradmatch $1.73\times 10^{-1}$ $1.48\times 10^{-1}$ $38\times 10^{-1} 1.1$$16\times 10^{-1}$ $1.03\times 10^{-1}$ $8.76\times 10^{-2}$
adacore $1.98\times 10^{-1}$ $1.58\times 10^{-1}$ $37\times 10^{-1} 1.1$$10\times 10^{-1}$ $1.02\times 10^{-1}$ $8.76\times 10^{-2}$
el2n $1.66\times 10^{-1}$ $1.48\times 10^{-1}$ $32\times 10^{-1} 1.1$$13\times 10^{-1}$ $1.03\times 10^{-1}$ $8.76\times 10^{-2}$
graNd $1.72\times 10^{-1}$ $1.47\times 10^{-1}$ $35\times 10^{-1} 1.1$$13\times 10^{-1}$ $1.03\times 10^{-1}$ $8.76\times 10^{-2}$